OpenStudy (anonymous):

Write the expression 81^-3/4 in radical notations Evaluate the radical expression.

OpenStudy (saifoo.khan):

\[\Huge 81^{-\frac34} \to \frac{1}{81^\frac34} \to \frac{1}{\sqrt[4]{81^3}}\]